The perimeter of a rectangle is 40 feet. The ratio of the width to the length is 2:3. Find the length and the width
Answer by london maths tutor(243)

Width : Length
2 : 3
Perimeter = 2* width + 2*length
Ratio becomes:
Width : Length
4 : 6
Total units = 4 + 6 = 10
Perimeter = 40
Therefore:
10 units ----- 40 ft
1 unit ------- 40/10 ft = 4 ft
2 units -------- 4 * 2 ft = 8 ft (width)
3 units ------- 4*3 ft = 12 ft (Length)
Answer: Width is 8 ft and Length is 12 ft.